Nguyen Inc. is considering the purchase of a new computer system (ICX) for $180,000. The system will require an additional $20,000 for installation. If the new computer is purchased it will replace an old system that has been fully depreciated. The new system will be depreciated over a period of 10 years using straight-line depreciation. If the ICX is purchased, the old system will be sold for $20,000. The ICX system, which has a useful life of 10 years, is expected to increase revenues by $33,000 per year over its useful life. Operating costs are expected to decrease by $1,000 per year over the life of the system. The firm is taxed at a 40 percent marginal rate. Round your answers to the nearest dollar.

  1. What net investment is required to acquire the ICX system and replace the old system?

    $  

  2. Compute the annual net cash flows associated with the purchase of the ICX system.

    $  
Answer #1

a)

Purchase price Add: Installation Less: old machine proceeds -$ 180,000 -$ 20,000 $ 12,000 =20000*(1-40%) Net Investment $ 188

b)

Revenue $ 34,000 =33000+1000 Less: Depreciation $ 20,000 =200000/10 Profit before tax $ 14,000 Less: Tax @ 40% $ 5,600 Net in

*Please rate thumbs up